OpenStudy (anonymous):

3(m+1.5) = 1.5 (2m+3) id or no solution

OpenStudy (anonymous):

identity

OpenStudy (saifoo.khan):

yes it does has identity. typo! >_<

OpenStudy (anonymous):

wats the answer

OpenStudy (anonymous):

3m+4.5=3m+4.5

OpenStudy (anonymous):

identity

OpenStudy (anonymous):

o ok thx :D

OpenStudy (anonymous):

i have 1 more 0.5b+4=2(b+2)

OpenStudy (saifoo.khan):

try this one of your own!

OpenStudy (anonymous):

i did and ppl said dat it had none

OpenStudy (anonymous):

0.5b+4=2b+4 0.5b=2b 0=1.5b b=0

OpenStudy (anonymous):

not any of the choices